Identify the greater number, wherever possible, in each of the following?
100^2 or 2^100

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The expansion of 100^2= 100 × 100 = 10000
The expansion of 2^100
2^10= 2 × 2 × 2 × 2 × 2 × 2 × 2 × 2 × 2 × 2 = 1024
Then,
2^100= 1024 × 1024 ×1024 × 1024 ×1024 × 1024 × 1024 × 1024 × 1024 × 1024 =
Clearly,
100^2< 2^100
Hence 2^100is the greater number.

Thanks